# Embodying Control in Soft Multistable Grippers from Morphofunctional Co-design

### v2

@@ A.6 Geometry Optimization
A.6.1 Inverse problem $\rightarrow$ Position Optimization
The final results for the position optimization can be observed in Figure S11, where the five target positions and the final shape predicted by the model are shown. The final geometry and objective function of the optimization algorithm are summarized in Table S3 and S4, where we can observe that for targets 1,2 and 5, the desired position is attained with almost perfect precision.
@@ A.6.1 Inverse problem $\rightarrow$ Position Optimization
A.6.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
We further expand our model to include stiffness optimization in our objective function. By incorporating finger stiffness, grasping force is increased by changing the geometric parameters. Results for the optimization algorithm, where the same number of segments as in Table S3 and S4 is maintained, are shown in Table S5 and S6.
@@ A.6.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
A.6.3 Inverse problem $\rightarrow$ Position + Stiffness Optimization for multiple states
We design the multistable actuator by targeting three different tip positions. These targets are intended to be reached by activating 2,4 and 6 units. The following objective function is minimized:
$\begin{split}\min_{H_{i},t,\text{U}^{i}_{L},\text{U}^{i}_{\text{sep}}}&\,w_{1}\left(\text{T1}_{xy}-\text{Tip}_{\text{dis 1}}\right)^{2}+w_{2}\left(\text{T2}_{xy}-\text{Tip}_{\text{dis 2}}\right)^{2}\\ &+w_{3}\left(\text{T3}_{xy}-\text{Tip}_{\text{dis 3}}\right)^{2}+w_{4}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}\end{split}$ (S.21)
s.t. $\quad H_{i+1}\geq H_{i},\quad i=1,\dots,N-1,$
$\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N,$
$\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N.$
Table S7: Experimental measurements for the multistable actuator.
Target Coordinates Model Experiments
State 1 ($T_{1}$) [54 , 40] [52.95 , 32.82] [ 50.2, 30.6]
State 2 ($T_{2}$) [40 , 45] [40.84 , 45.9] [ 38.2, 44.2]
State 3 ($T_{3}$) [30 , 45] [33.81 , 47.35] [ 35.2, 46.1]
Where $w_{1}=w_{2}=w_{3}=w_{4}=1/4$ and $T_{1}$ to $T_{3}$ represent the target positions for the gripper as shown in Table S7. Results for each of the targets can be seen in Figure S12.
Figure S12: Inverse optimization results for multistable dome phalanx actuator. a) Lattice model results for $T_{1}$(i), $T_{2}$(ii) and $T_{3}$(iii). b) Experimental results for $T_{1}$(i), $T_{2}$(ii) and $T_{3}$(iii).
A.6.4 Inverse problem $\rightarrow$ Bistable + Mestastable units
Our task planning application uses our dynamic model to calculate the resetting time of the metastable units. We fix our actuator geometry by selecting six segments with four bistable and two metastable units. The selected geometry parameters are $H_{1}=5$ mm, $H_{2}=5$ mm and $H_{3}=5$ mm, $\text{U}_{\text{L}}=6.5$ mm and $\text{U}_{\text{sep}}=1$ mm, $t=0.8$ mm, and the last two-unit heights ($H_{5}$ and $H_{6}$) is designed to maximize the resetting time under the two different pressure profiles show in Figure S13a. Both dome heights are constrained to be the same ($H_{5}=H_{6}$), and the initial bistable set point can be observed in Figure S13b. The full dynamic response predicted by the model is shown in Figure S13c, where we can observe the three different snap-through instability and two snap-backs in the two simulated cases. A restriction is imposed over these units to guarantee that they were metastable, by using the state map reported Figure S1b. As expected, to maximize the resetting time at two different conditions, the algorithm would converge to the highest possible metastable unit (see Figure S1b).
@@ A.7.2 Inverse problem experimental validation
Five geometries are tested by activating all bistable unit cells and measuring the final tip position. The final position of the finger is measured by using image processing. Results can be observed in Figure S15a (i-v) for NinjaFlex 85A and Figure S15b (i-iii) for Cheetah 95A, and the target coordinates, model prediction, and experimental measurements are reported in Table S9
@@ A.7.3 Experimental validation stiffness optimization
The stiffness of each DPF is measured using an Instron universal testing machine and applying a perpendicular force to the tip of the actuator (see Figure S15b and Movie 5). Linear behavior is observed in the neighborhood of each stable state (0, 2, 4, and 5 active units in Figure S15b), with a different stiffness for every state. Finally, the inverse design problem results are tested by comparing the stiffness of two geometries with the same target position but two distinct objective functions (Equation 6 and Equation 8).
@@ A.7.5 Carrying Capacity
Payload capacity and object grasping tests were conducted to evaluate the performance of our DPR, demonstrate its grasping capabilities, and determine the maximum carrying weight of the system. The designed global states enable adjustments in output force and position based on the number of inverted domes (see Movie 6). This adaptability allows the DPR to grasp objects of various sizes and topologies by removing the actuation pressure. Figure S18 illustrates the range of objects successfully grasped by the DPR using only the energy stored in the structure due to dome inversion. As depicted, diverse object types are effectively grasped, highlighting the system’s versatility. Additionally, to characterize the payload capacity, weights were incrementally added to the object shown in Figure S18 (i) until the structure could no longer grasp it. A maximum load of 738 g was recorded, resulting in a maximum load-to-weight ratio of 10.25 (See Movie 6).
@@ A.7.6 Scalability
As the mechanical response of the DPF is derived from the dome geometry and its stability, the DPF and robot can be scaled to increase to large dimensions just by considering the energy stored during the dome inversion. Following the analysis by Seffen and Vidoli [58], we can derive different additional relations for scaling, given the energy required for inverting a single dome. This energy is proportional to $\frac{Et^{3}}{1-\nu^{2}}\left(\frac{H}{R}\right)^{2}$ where E, $\nu$, t, and R are the elastic modulus, Poisson’s ratio, thickness, height, and radius of the domes respectively.
We can utilize the dome shallowness $\frac{H}{R}$ and the curvature-to-thickness ratio $\frac{t}{R}$ to either scale up or down the DPF geometry. To test this, we 3D printed two different geometries with the same energetics but a 2x difference.

### v3

@@ A.6 Geometry Optimization
A.6.1 Design space for inverse design
We explored the design space of our model by calculating the tip displacement, curvature, and stiffness as a function of the input parameters. Given the width range of parameters, we utilize a Latin Hypercube sampling technique to uniformly sample all of the parameter space ($H_{i}$, $\text{U}^{i}_{\text{sep}}$, $\text{U}^{i}_{\text{L}}$, $t$ and $t_{\text{lim}}$) of the DPF to observe its response after all the units are activated. 1500 samples are taken to guarantee the full exploration of the design space. Analysis for different numbers of units (2 - 7 units) is performed, and the design variables are kept in the ranges listed in Table S3.
min max
$H_{i}$ 3.0 5.0
$\text{U}^{i}_{\text{sep}}$ 1.0 5.0
$\text{U}^{i}_{\text{L}}$ 5.5 10.0
$t$ 0.5 1.0
$t_{\text{lim}}$ 1.0 2.0
Table S3: Optimization variables ranges.
Figure S11: Exploration of the design space. We use our spring lattice model to characterize the deformation modes of structures made of different numbers of units with varying input parameters (Distributed uniformly). a) Tip position design space. b) Curvature design space. c) Stiffness design space.
A.6.2 Inverse problem $\rightarrow$ Position Optimization
Given a target tip coordinate [x,y] ($\text{Target}_{xy}$), the inverse problem objective function can be written as:
$\min_{H_{i},t,\text{U}^{i}_{L},\text{U}^{i}_{\text{sep}}}\left(\text{Target}_{xy}-\text{Tip}_{\text{dis}}\right)^{2}$ (S.30)
[…]
The final results for the position optimization can be observed in Figure S12, where the five target positions and the final shape predicted by the model are shown. The final geometry and objective function of the optimization algorithm are summarized in Table S4 and S5, where we can observe that for targets 1,2 and 5, the desired position is attained with almost perfect precision.
@@ A.6.2 Inverse problem $\rightarrow$ Position Optimization
A.6.3 Inverse problem $\rightarrow$ Position + Stiffness Optimization
We further expand our model to include stiffness optimization in our objective function. The finger’s stiffness is calculated by using the dynamic model described in section A.4, by perturbing the structure from a programmed stable configuration (i.e., set point). We determine the system’s total internal force by solving the dynamic equation:
$[M]\ddot{x_{i}}+F_{d}^{int}(x_{ij},\dot{x_{i}},\dot{x_{j}})+F_{d}^{iso}(\dot{x_{i}},\dot{x_{j}})+F_{\text{in}}=F_{ext}(x_{i,j},t)$ (S.31)
$F_{\text{in}}=\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)$
where $F_{\text{ext}}(x_{i,j},t)$ is a follower force applied at the tip of the DPF (see Figure 4d) to perturb the DPF from its equilibrium position. The gradient of the internal forces in the structure, $F_{\text{in}}$, is obtained through the analytical expression described in Section 2.2. The force is applied at a constant load rate of 5 mm/min, ensuring a quasi-static regime. The stiffness of the DPF in the desired state is calculated by determining the internal force ($F_{\text{in}}$) and the corresponding tip displacement over time (see Figure 4). A linear fit is then applied to the resulting data. Using this methodology, the DPF’s geometry can be optimized to achieve a specific target position with maximum stiffness in the vicinity of that stable state (see Figure 4e).
By incorporating finger stiffness, grasping force is increased by changing the geometric parameters. Results for the optimization algorithm, where the same number of segments as in Table S4 and S5 is maintained, are shown in Table S6 and S7.
@@ A.7.2 Inverse problem experimental validation
Five geometries are tested by activating all bistable unit cells and measuring the final tip position. The final position of the finger is measured by using image processing. Results can be observed in Figure S14a (i-v) for NinjaFlex 85A and Figure S14b (i-iii) for Cheetah 95A, and the target coordinates, model prediction, and experimental measurements are reported in Table S9.
@@ A.7.3 Experimental validation stiffness optimization
The stiffness of each DPF is measured using an Instron universal testing machine and applying a perpendicular force to the tip of the actuator (see Figure S14b). Linear behavior is observed in the neighborhood of each stable state (0, 2, 4, and 5 active units in Figure S14b), with a different stiffness for every state. Finally, the inverse design problem results are tested by comparing the stiffness of two geometries with the same target position but two distinct objective functions (Equation S.30 and Equation 6).
@@ A.7.5 Carrying Capacity
Payload capacity and object grasping tests were conducted to evaluate the performance of our DPR, demonstrate its grasping capabilities, and determine the maximum carrying weight of the system. The designed global states enable adjustments in output force and position based on the number of inverted domes (see Movie 5). This adaptability allows the DPR to grasp objects of various sizes and topologies by removing the actuation pressure. Figure S17 illustrates the range of objects successfully grasped by the DPR using only the energy stored in the structure due to dome inversion. As depicted, diverse object types are effectively grasped, highlighting the system’s versatility. Additionally, to characterize the payload capacity, weights were incrementally added to the object shown in Figure S17 (i) until the structure could no longer grasp it. A maximum load of 738 g was recorded, resulting in a maximum load-to-weight ratio of 10.25 (See Movie 5).
@@ A.7.6 Scalability
As the mechanical response of the DPF is derived from the dome geometry and its stability, the DPF and robot can be scaled to increase to large dimensions just by considering the energy stored during the dome inversion. Following the analysis by Seffen and Vidoli [58], we can derive different additional relations for scaling, given the energy required for inverting a single dome. This energy is proportional to $\frac{Et^{3}}{1-\nu^{2}}\left(\frac{H}{R}\right)^{2}$ where E, $\nu$, t, and R are the elastic modulus, Poisson’s ratio, thickness, height, and radius of the domes respectively. We can utilize the dome shallowness $\frac{H}{R}$ and the curvature-to-thickness ratio $\frac{t}{R}$ to either scale up or down the DPF geometry. It should be mentioned that our dome unit bistable behavior can be geometrically scaled; however, for metastable units, it is necessary to consider the effect of the viscoelastic response on a larger system as this response is a combination of the geometry and material response. To test this, we 3D printed two different geometries with the same energetics but a 2x difference.
@@
A.8 Dynamic response for Pick and Place Application
For our pick and place, we create a pressure profile that accurately matches the experimental measurements for our DPF (see Figure S19a). At the same time, the optimization algorithm is utilized to determine the dome heights of the last two units and the remaining geometric parameters ($t$).
Figure S20: Programmed Pick and Place application for the Dome Phalanx Gripper a) Close up of commanded response when picking a tennis ball. b) Bistable + Metastable finger behavior: Strain energy and pressure vs. time response. c) Snapshots of embodied pick-and place tasks: A single pressure input is given for the robot to pick the object (Commanded response) and then release it depending on the viscoelastic material response (Programmed response)(see Movie 7).
Table S10: Pick and Place DPF final geometry
H1 4.0
H2 4.5
H3 5.0
H4 5.0
H5 3.0
H6 3.0
Unit S 1.0
Unit L 6.0
t 0.83
$t_{\text{lim}}$ 1.0
UC 15.00
Given this, a DPG architecture is built that utilizes the dynamic programmed response (see Figure S20a) to perform a simple pick-and-place application after the pressure is released. This behavior is achieved as the metastable domes reset due to visco-elastic effects, showing a snap-back phenomenon on a specific time scale (see Figure S20). It should be mentioned that the metastability can be utilized in either stable state, meaning this embodied task can be performed for different gripper apertures (see Movie 6).
A.8.1 Embodied Classification Task
We leverage the multistability of our DPF to develop a multistable gripper (DPG) capable of achieving multiple aperture configurations. This enables the robot to approximate and adapt to various object sizes. Representative object sizes, corresponding target positions, and optimization results are presented in Table S11, while a visual comparison with the 3D-printed prototypes is shown in Figure S21b. Due to the gripper’s symmetry, the aperture is modeled considering only two fingers. By targeting a specific set of aperture values, we encode object size information directly into the gripper’s geometry and design. This allows for discrimination between objects with dimensions near the predefined targets. To decode this morphological information, we integrate a contact sensor into one of the finger units (Figure S21c), which detects when the gripper deviates from a stable state. We use a commercial RP-C10-ST pressure sensor (0.2–2 kg), although other sensors capable of detecting perturbations—such as curvature or strain sensors—can also be employed. The sensor threshold is calibrated by observing the sensor reading when the dome behind it is inverted, guaranteeing that all the stable configurations would not be detected as a contact. By utilizing the feedback from the contact sensor and the robot’s morphology, we can create a classification loop (see Movie 7) that can classify between different object sizes and weights.
Table S11: Object size and target coordinate for multistable DPR
Object size [mm] Model
35 35.02
45 45.07
60 60.023
70 73.15
Figure S21: Dome Phalanx Gripper lattice design. a) Lattice model results for the four optimized stable states. b) 3D printed DPG on its four different programmed stable states. Black disks represent the designed object size. c) Contact sensor position placement to detect perturbations from the pre-programmed stable state.
Results for the optimization algorithm for different object sizes can be observed in Table S12.
Table S12: Object classification optimization results
H1 4.82
H2 4.52
H3 4.15
H4 3.18
H5 2.5
Unit S 1.82
Unit L 8.57
t 0.83
$t_{\text{lim}}$ 1.05
UC 15.00
Base L 32.4
$\theta_{B}$ 28
A.9 Dome Phalanx Walker
The dome phalanx walker (Figure S22a) is designed by combining six legs, each composed of different arrangements of dome units (see Figure S22a). The front and back legs are angled at 30, inspired by leg arrangements observed in insects [59], to enhance both stability and locomotion speed. Each leg has two functional zones (Purple and green zone in Figure S22c), each composed of a combination of different dome unit geometries which can exhibit monostable, metastable, and bistable behaviors. Each leg exhibits two stable states (Figure S22b), enabling the robot to support greater loads while remaining compatible with monolithic 3D printing in its initial flat configuration. To study the leg’s motion and design the resetting times of the metastable units, we employ our lattice model to generate a discrete representation of the leg (see Figure S22d). Our central modeling assumption is that units within different planes and zones do not interact, allowing for independent design of each zone while maintaining the model’s simplicity. Using this approach, we determine both the metastable units’ resetting times and each dome unit’s design parameters (Figure S22c) to program the timing of each locomotion phase. Additionally, the model allows us to track the tip displacement in the y-z plane over time (Figure S22e), revealing how metastability amplifies leg movement through the time delays introduced by the viscoelastic response.
Figure S22: Dome Phalanx Walker design with lattice model. a) DPW architecture. b) Flat state of the walker (State 1) and arc-like state (State 2) by activating the bistable units. c) The DPW leg and its dynamic behavior are divided by zones and units. The metastable action unit exhibits a higher resetting time than the metastable units in the purple region.
The design unit (see Figure S22c) is tuned individually for each leg, enabling versatility and programmability in the walker’s behavior. We employ this unit to achieve directional walking by controlling only the actuation pressure (see Figure S23a and b). This is accomplished by introducing a design unit with metastable behavior in the front legs, characterized by a longer resetting time and a higher inversion pressure than the metastable unit used in zone 1 (purple zone).This configuration ensures that the front-leg units do not prematurely revert to their initial state, as actuation pressure is applied before their resetting time elapses. This creates a sustained asymmetry in leg behavior, driving directional movement until the pressure is released or increased on the opposite phase (see Figure S23). It is important to note that the dome geometry directly affects the resetting time—longer resetting times are associated with higher inversion pressures. Accordingly, we designed the front-leg units for a longer resetting time, resulting in higher inversion pressure.
Figure S23: Dome Phalanx Walker programmed response and pressure measurements a) Snapshots of a straight walking direction given a uniform alternating pressure phase. b) Snapshots of the turning maneuver result from an asymmetry created by the pressure increase on the first phase cycle.